OpenStudy (anonymous):

Find t in the equation: 1/2=(0.87)^t

OpenStudy (aravindg):

simple it asks for what value of t , .87^t gives 1/2

OpenStudy (aravindg):

for t=5 we get 4.9 approx .5=1/2

OpenStudy (aravindg):

its .49

OpenStudy (auctoratrox):

take the natural log of both sides, then divide both sides by ln(0.87) to get t = ln(1/2)/ln(0.87)

OpenStudy (anonymous):

I'm new to this kind of maths, I was wondering what the In means in your final answer?

OpenStudy (auctoratrox):

it means the natural log. the natural log is the inverse of the exponential function, y = e^x.

OpenStudy (anonymous):

Thankyou!

OpenStudy (lgbasallote):

it seems you have not discussed ln yet? you can also change it to logarithmic form (which im sure you are familiar with) \[a^x = b \rightarrow \log_a b = x\] so in this case, we change it to logarithmic form... \[\Large\log_{\frac{1}{2}} (0.87) = t\] which can then be solved by calculator. "ln" is commonly done in calculus so i assume that's why @aravindg used it